OpenStudy (anonymous):

a model rocket is launched from a raised plat form at aspeed of 192 ft per second its height is given by h(t)= -16t^2+192t+20 (t=secs after launch) about how many seconds does the object reach its maximum height? A) 20 B) 6 C) 3 D) 8

OpenStudy (bahrom7893):

Well how do you find the maximum?

OpenStudy (anonymous):

thats what im confused about

OpenStudy (bahrom7893):

Take the derivative and set it equal to 0, then solve for t

OpenStudy (bahrom7893):



OpenStudy (anonymous):

i get 6

OpenStudy (bahrom7893):

Don't guess. How did you end up with 6?

OpenStudy (anonymous):

set all the t to 0 right?

OpenStudy (bahrom7893):

Wait what class is this for? Calculus?

OpenStudy (anonymous):

college alg

OpenStudy (bahrom7893):

do you know how to take derivatives?

OpenStudy (anonymous):

not really

OpenStudy (bahrom7893):

Well, are you supposed to know how to take derivatives in that class?

OpenStudy (anonymous):

prof never said anything about it

OpenStudy (bahrom7893):

Okay, so since you guys don't do derivatives, I guess we can do this another way. Just find the vertex of that parabola.

OpenStudy (anonymous):

ugh i dont get it :/

OpenStudy (bahrom7893):

Get that equation into the form: y = a(x – h)^2 + k

OpenStudy (anonymous):

idk how

OpenStudy (bahrom7893):

-16t^2+192t+20 is in the form at^2+bt+c We need to get it in a form: y=a(x-h)^2+k

OpenStudy (bahrom7893):

First rewrite the equation as: -(16t^2-192t-20) = -( (4t)^2 - 2*4*24t - 20 )

OpenStudy (bahrom7893):

then complete the square to get it in the form we need it in

OpenStudy (anonymous):

-20

OpenStudy (anonymous):

after i solve it

OpenStudy (bahrom7893):

No don't solve it. Get it into the form that i was talking about, and that will give you your (h; k). h will be your answer

OpenStudy (anonymous):

i dont know what variable represents what number

OpenStudy (bahrom7893):

Ok man, -32t+192=0. Solve for t, that will be your answer